OpenStudy (anonymous):

The length of a rectangle is 10cm and the width is 6sm. Each side is enlarged by x cm. The enclosed area of the new rectangle is 480 cm. Whats the new dimensions of the ractangle, and how?

OpenStudy (anonymous):

(10+x)(6+x)=480

OpenStudy (anonymous):

60 + 16x +x^2=480

OpenStudy (anonymous):

I got -30 by 14 which is impossible

OpenStudy (anonymous):

x^2+16x=420

OpenStudy (anonymous):

i used the quadratic formula

OpenStudy (anonymous):

x^2+16x-420 (x-14)(x+30)=0 x=14 or -30 it can be negative so x=14

OpenStudy (anonymous):

and then 10+14 and 6+14

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks!